# Compression for Adaptive Edge LLMs

### v2

@@ 3.3 Masked LoRA Fine-tuning
We conduct a LoRA-based (hu2022lora) recovery fine-tuning (FT) on the sorted model. Unlike previous work (wang2025svdllm; wang2025svd) that fine-tunes the pruned model, we fine-tune the un-pruned sorted model in one shot, as shown in Figure 2. We derive the layer-wise pruning rates $r_{l}$ using Block Influence (BI) scores (lin2025modegpt; men2024shortgpt) for all pre-determined global pruning rates $P=[P_{15},P_{20},$...], such as $P_{15}=[r^{P_{15}}_{1},r^{P_{15}}_{2},...,r^{P_{15}}_{L}]$. The BI score is defined as $s=1-\mathbb{E}\frac{\mathbf{x}_{l}^{\top}\mathbf{y}_{l}}{\|\mathbf{x}_{l}\|_{2}\|\mathbf{y}_{l}\|_{2}}$, where the $x_{l}$ and $y_{l}$ represent the input and output of the $l^{\mathrm{th}}$ layer, respectively. During fine-tuning, we randomly draw a pruning rate $P_{t}\sim P$ at time step $t$ to mask out the pruned channels. We follow the prior work (wang2025svdllm) to perform instruction tuning on the Alpaca dataset (taori2023stanford) for five epochs. The entire fine-tuning process is conducted on a single cloud GPU. Our sorted model provides configurable pruning rates on the device after one-shot masked fine-tuning. We note that our fine-tuning inherently supports downstream tasks for any application, such as summarization and question-answering datasets.
@@ 3.4 Quantization and On-device Adaptive Pruning
where $s=\text{\footnotesize{Max}}\big(\left\lvert\boldsymbol{W}_{(i,g)}\right\rvert\big)/(2^{N-1}-1)$ is the scaling factor (i.e., quantization step). For the quantization-aware SVD decomposition, we fuse the eigenvalues of the $i^{\mathrm{th}}$ column to the group of weight factor such that $\sigma_{i}\boldsymbol{W}_{(i,g)}$. We set $N=4$ and group size $128$ and adapt GPTQ (frantar2023gptq) to quantize our models. We fuse Hadamard matrices into the weights and apply 4-bit quantization to the embedding and output layers. The parameters of normalization layers are fused to the weights before applying quantization. After deploying the quantized model, we prune the channels on the device by reducing the intermediate dimension $D_{\mathrm{int}}$ in the MLP layer, head dimension $D_{\mathrm{hd}}$ in the MHSA layer, and both the state dimension $D_{\mathrm{s}}$ and head dimension $D_{\mathrm{hd}}$ in the Mamba layer. We keep the hidden state dimension $D_{\mathrm{h}}$ the same across all pruned models. For INT4 weights, we unpack them online, prune channels, and repackage them into INT32 for the kernel.
@@
Prun.%
Prun. Method
+FT
Llama-2-7B
Llama-3.1-8B
Qwen-2.5-7B
Bamba-v2-9B
Nemotron-H-8B
Mamba2-8B
@@
Table 3: (Compared with PTQ.) We benchmark all weight-only PTQ methods without fine-tuning and pruning. represents the FP16 embeddings and output layers as per the official implementation. denotes the GPTQ (frantar2023gptq) implemented on all models as an additional baseline.
PTQ Method
W-bit
Llama-2-7B
Llama-3.1-8B
Qwen-2.5-7B
Bamba-v2-9B
Nemotron-H-8B
Mamba2-8B
@@
UniQL (Ours)
4111The embedding and output layers use FP16 according to the official implementation. 68.2% 72.9% 72.0% 74.8% 74.9% 69.3%
@@ 4.1 Setups
We experiment with Transformers Llama-2-7B (touvron2023llama2), Llama-3.1-8B (meta2024llama), Qwen-2.5-7B (hui2024qwen2), hybrid models Nemotron-H-8B (blakeman2025nemotron), Bamba-9B-v2 (bamba2025v2), and the SSM model Mamba-2-8B (dao2024transformers). $\diamond$, $\dagger$, and $\ddagger$ denote Transformers, hybrid and SSMs, respectively. FT, PTQ, and W-bit stand for fine-tune, post-training quantization, and the bit-width of weights. Prun. and R.size represent the pruning rate in percentage (%) and the reduction of model size ($\times$), respectively.
@@
We compare UniQL to cutting-edge model compression methods, MoDeGPT (lin2025modegpt) and SVD-LLM (wang2025svdllm). As MoDeGPT is not publicly available, we duplicate their method based on the paper and achieve similar accuracy to what was reported. We adapt SVD-LLM official implementation to experiment with Llama-2-7B, Llama-3.1-8B, and Qwen-2.5-7B for compression and quantize models. For MoDeGPT and SVD-LLM, we adhere to the hyper-parameters outlined in the papers.
@@
We adopt AWQ (lin2024awq) in TensorRT-MO [^1] The embedding and output layers use FP16 according to the official implementation. (TRT-AWQ) (nvidia2024trtmo; nvidia2023tensorrtllm) and HQQ (badri2023hqq) in TorchAO [^1] The embedding and output layers use FP16 according to the official implementation. (torchao) (TAO-HQQ) as our quantization baselines, both W4A16 libraries ready for PTQ. We evaluate UniQL in terms of model size, average accuracy on downstream tasks, and latency on A6000 and Nano 8G against TRT-AWQ and TAO-HQQ. In our experiments, we quantize the embedding and output (i.e., lm_head) layers to 4 bits, cutting memory usage in contrast to TRT-AWQ [^1] The embedding and output layers use FP16 according to the official implementation. and TAO-HQQ [^1] The embedding and output layers use FP16 according to the official implementation. which use FP16. We also adapt GPTQ (frantar2023gptq) for all models and also use 4-bit quantization on the embedding and output layers as an additional baseline.
@@
We evaluate UniQL on five zero-shot tasks with a batch size of 16, including HellaSwag (HellaSwag), PIQA (PIQADataset), ARC (arcDataset), and WinoGrande (WinoGrandeDataset) using LM-EVAL (eval-harness). The average of WinoGrande, PIQA, and ARC-easy (accuracy), and HellaSwag and ARC-challenge (length-normalized accuracy) is reported for experiments. More evaluations are placed in Appendix B.
@@
Our kernels are adapted from the 4-bit kernels (frantar2024marlin) and RoPE kernels (hsu2025ligerkernel). All computations are in BF16, except for correlation matrices for structured sorting and Cholesky decomposition for GPTQ are calculated in FP32. Detailed parameters are placed at Appendix G. The weight-sorting, masked fine-tuning, and quantization are computed on an A6000 GPU with 48GB memory in one-shot for enabling adaptive pruning on the device. We profile the latency on A6000 and Orin Nano 8GB, as our experimental cloud and edge platforms. We report the average latency of twenty profiles after five warmup runs.
Table 4: (One-pass adaptive pruning.) We evaluate UniQL in one run across pruning rates. R.size stands for reduction of model size ($\times$). UniQL enables all pruning rates in single pass. The symbols $\diamond$, $\dagger$, and $\ddagger$ denote Transformers, Mamba-Transformer hybrids, and Mamba models, respectively. represents the FP16 embeddings and output layers as per the official implementation. We apply GPTQ (frantar2023gptq) on MoDeGPT (lin2025modegpt) denoted as .
Method
one
pass
+FT
W
bit
Prun.
R.size
($\times$)
Llama-2
7B
Llama-3.1
8B
Qwen-2.5
7B
Bamba-v2
9B
Nemotron-H
8B
Mamba2
8B
@@ 4.3 Compression Time
In Table 6, we compare the compression time of UniQL against MoDeGPT (lin2025modegpt), noted for being training-free, and SVD-LLM (wang2025svdllm), which involves fine-tuning, both state-of-the-art algorithms for Transformer compression. Our matrix decomposition is $22\times$ (0h19m vs. 7h03m) faster than MoDeGPT and $1.8\times$ (0h19m vs. 0h35m) faster than SVD-LLM, as UniQL avoids pseudo-inverse and SVD decomposition for large MLP weight matrices. With masked fine-tuning (FT), UniQL remains quicker (6h59m) than both MoDeGPT (7h03m) and SVD-LLM (15h57m). MoDeGPT suffers from the high computation cost of performing pseudo-inverse on large weight matrices. SVD-LLM splits weights into two successive layers, $\mathbf{U}$ and $\mathbf{V}$, and carries out independent fine-tuning for each, leading to a longer fine-tuning duration. Lastly, post-training quantization (PTQ) takes an extra forty minutes. Our compression algorithm is one-time $O(1)$ with respect to the number of compression rates, compared to $O(n)$ for MoDeGPT and SVD-LLM.
@@ 4.4 Model Size and Latency Profiling
We evaluate the model size and latency of UniQL and compare them with AWQ (lin2024awq) from TensorRT-MO11footnotemark: 1 (nvidia2024trtmo; nvidia2023tensorrtllm) (TRT-AWQ) and HQQ (badri2023hqq) in TorchAO11footnotemark: 1 (torchao) (TAO-HQQ). Both libraries are weight-only (i.e., W4A16) quantization frameworks in production. We show the model size in Table 6. UniQL quantizes models in head-to-toe fashion, i.e., from embeddings, backbone layers, and output heads all at 4-bit, resulting in smaller model size (4.1GB vs. 5.7GB) compared to TRT-AWQ and TAO-HQQ with minimal accuracy drops. Our model enables all compression rates and on-device structured pruning, providing an elastic $3.9\times$$5.7\times$ memory reductions. We profile the time-per-output-token (TPOT, i.e., generation) and time-to-last-token (TTLT, i.e., prefilling and generation) on A6000 and Nano 8G, and show $2.7\times$$3.4\times$ throughput improvements in generation, outperforming TRT-AWQ and TAO-HQQ, as shown in Table 8 and Table 8. On the Nano 8G, our model is $1.7\times$ faster than TAO-HQQ in TPOT. By pruning 35% of weights in our 4-bit model, our models generate $2.1\times$ faster than TAO-HQQ.
@@ Acknowledgments
This work was supported in part by the ONR Minerva program, NSF CCF Grant No. 2107085, iMAGiNE - the Intelligent Machine Engineering Consortium at UT Austin, UT Cockrell School of Engineering Doctoral Fellowships, NSF CAREER Grant No. 2339084, Nvidia research gift, and Taiwan’s NSTC Grant No. 111-2221-E-A49-148-MY3.
@@ B.1 Comparison with Additional Baselines
we include more common baselines in Table 11, with all methods evaluated in FP16 to ensure a fair and controlled setting. We follow the experimental setup used in MoDeGPT [lin2025modegpt] and append our Llama-3.1-8B results at the 25% pruning rate to those reported in their work. The numbers, except for UniQL, are transcribed from MoDeGPT. The results show that UniQL consistently outperforms prior pruning-based methods at the 25% compression level, and UniQL-ft further boosts accuracy, achieving the strongest performance across all evaluated tasks.
@@ B.2 Evaluation on the MMLU Dataset
We test Llama-3.1-8B, Qwen-2.5-7B, and Nemotron-H-8B, and report five-shot accuracy on the MMLU dataset [hendrycks2020measuring] with a batch size of eight. The MMLU dataset is a large multitasking dataset, covering 57 subjects of varying difficulty. Our pruned models maintain competitive accuracy on the challenging dataset and outperform MoDeGPT [lin2025modegpt] and SVD-LLM [wang2025svdllm]. Nemotron-H, the Mamba-Transformer hybrid model, shows a large accuracy drop, but recovered with our low-cost masked fine-tuning. We compare our method with AWQ [lin2024awq] implemented in the TensorRT framework [nvidia2023tensorrtllm, nvidia2024trtmo] (TRT-AWQ) and HQQ [badri2023hqq] in TorchAO [torchao] (TAO-HQQ). Our 4-bit models perform comparably to these state-of-the-art PTQ frameworks while offering broader model support.
Table 12: (Five-shot MMLU.) We compare UniQL against baselines under different settings. represent the FP16 embeddings and output layers as per the official implementation. denotes the GPTQ [frantar2023gptq] implemented on all models as an additional baseline.
Method
+FT
+PTQ
W-bit
Prun. p%
Llama-3.1-8B
Qwen-2.5-7B
Nemotron-H-8B
@@ B.3 Evaluation on Coding Tasks
We present the evaluation results on the MBPP+ [austin2021program, eval-harness] coding benchmark in Table 13. We apply UniQL to Llama-3.1-8B-Instruct and compare its performance against SVD-LLM [wang2025svdllm] and MoDeGPT [lin2025modegpt] under various pruning ratios and bit-width configurations. The MBPP+ results obtained under batch size 1 and 0-shot settings following common practice. These results demonstrate UniQL’s ability to maintain competitive performance compared to SVD-LLM and MoDeGPT while reducing model size.

### v3

@@ 3.3 Masked LoRA Fine-tuning
We conduct a LoRA-based (Hu et al., 2022) recovery fine-tuning (FT) on the sorted model. Unlike previous work (Wang et al., 2025b, a) that fine-tunes the pruned model, we fine-tune the un-pruned sorted model in one shot, as shown in Figure 2. We derive the layer-wise pruning rates $r_{l}$ using Block Influence (BI) scores (Lin et al., 2025; Men et al., 2024) for all pre-determined global pruning rates $P=[P_{15},P_{20},$...], such as $P_{15}=[r^{P_{15}}_{1},r^{P_{15}}_{2},...,r^{P_{15}}_{L}]$. The BI score is defined as $s=1-\mathbb{E}\frac{\mathbf{x}_{l}^{\top}\mathbf{y}_{l}}{\|\mathbf{x}_{l}\|_{2}\|\mathbf{y}_{l}\|_{2}}$, where the $x_{l}$ and $y_{l}$ represent the input and output of the $l^{\mathrm{th}}$ layer, respectively. During fine-tuning, we randomly draw a pruning rate $P_{t}\sim P$ at time step $t$ to mask out the pruned channels. We follow the prior work (Wang et al., 2025b) to perform instruction tuning on the Alpaca dataset (Taori et al., 2023) for five epochs. The entire fine-tuning process is conducted on a single cloud GPU. Our sorted model provides configurable pruning rates on the device after one-shot masked fine-tuning. We note that our fine-tuning inherently supports downstream tasks for any application, such as summarization and question-answering datasets.
@@ 3.4 Quantization and On-device Adaptive Pruning
where $s=\text{\footnotesize{Max}}\big(\left\lvert\boldsymbol{W}_{(i,g)}\right\rvert\big)/(2^{N-1}-1)$ is the scaling factor (i.e., quantization step). For the quantization-aware SVD decomposition, we fuse the eigenvalues of the $i^{\mathrm{th}}$ column to the group of weight factor such that $\sigma_{i}\boldsymbol{W}_{(i,g)}$. We set $N=4$ and group size $128$ and adapt GPTQ (Frantar et al., 2023) to quantize our models. We fuse Hadamard matrices into the weights and apply 4-bit quantization to the embedding and output layers. The parameters of normalization layers are fused to the weights before applying quantization. After deploying the quantized model, we prune the channels on the device by reducing the intermediate dimension $D_{\mathrm{int}}$ in the MLP layer, head dimension $D_{\mathrm{hd}}$ in the MHSA layer, and both the state dimension $D_{\mathrm{s}}$ and head dimension $D_{\mathrm{hd}}$ in the Mamba layer. We keep the hidden state dimension $D_{\mathrm{h}}$ the same across all pruned models. For INT4 weights, we unpack them online, prune channels, and repackage them into INT32 for the kernel.
@@
Prun.% Prun. Method +FT Llama-2-7B Llama-3.1-8B Qwen-2.5-7B Bamba-v2-9B Nemotron-H-8B Mamba2-8B
@@
Table 3: (Compared with PTQ.) We benchmark all weight-only PTQ methods without fine-tuning and pruning. represents the FP16 embeddings and output layers as per the official implementation. denotes the GPTQ (Frantar et al., 2023) implemented on all models as an additional baseline.
PTQ Method W-bit Llama-2-7B Llama-3.1-8B Qwen-2.5-7B Bamba-v2-9B Nemotron-H-8B Mamba2-8B
@@
UniQL (Ours) 4111The embedding and output layers use FP16 according to the official implementation. 68.2% 72.9% 72.0% 74.8% 74.9% 69.3%
@@ 4.1 Setups
We experiment with Transformers Llama-2-7B (Touvron et al., 2023), Llama-3.1-8B (Meta, 2024), Qwen-2.5-7B (Hui et al., 2024), hybrid models Nemotron-H-8B (Blakeman et al., 2025), Bamba-9B-v2 (IBM, 2025), and the SSM model Mamba-2-8B (Dao and Gu, 2024). $\diamond$, $\dagger$, and $\ddagger$ denote Transformers, hybrid and SSMs, respectively. FT, PTQ, and W-bit stand for fine-tune, post-training quantization, and the bit-width of weights. Prun. and R.size represent the pruning rate in percentage (%) and the reduction of model size ($\times$), respectively.
@@
We compare UniQL to cutting-edge model compression methods, MoDeGPT (Lin et al., 2025) and SVD-LLM (Wang et al., 2025b). As MoDeGPT is not publicly available, we duplicate their method based on the paper and achieve similar accuracy to what was reported. We adapt SVD-LLM official implementation to experiment with Llama-2-7B, Llama-3.1-8B, and Qwen-2.5-7B for compression and quantize models. For MoDeGPT and SVD-LLM, we adhere to the hyper-parameters outlined in the papers.
@@
We adopt AWQ (Lin et al., 2024a) in TensorRT-MO [^1] The embedding and output layers use FP16 according to the official implementation. (TRT-AWQ) (NVIDIA, 2024, 2023) and HQQ (Badri and Shaji, 2023) in TorchAO [^1] The embedding and output layers use FP16 according to the official implementation. (torchao, 2024) (TAO-HQQ) as our quantization baselines, both W4A16 libraries ready for PTQ. We evaluate UniQL in terms of model size, average accuracy on downstream tasks, and latency on A6000 and Nano 8G against TRT-AWQ and TAO-HQQ. In our experiments, we quantize the embedding and output (i.e., lm_head) layers to 4 bits, cutting memory usage in contrast to TRT-AWQ [^1] The embedding and output layers use FP16 according to the official implementation. and TAO-HQQ [^1] The embedding and output layers use FP16 according to the official implementation. which use FP16. We also adapt GPTQ (Frantar et al., 2023) for all models and also use 4-bit quantization on the embedding and output layers as an additional baseline.
@@
We evaluate UniQL on five zero-shot tasks with a batch size of 16, including HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), ARC (Clark et al., 2018), and WinoGrande (Sakaguchi et al., 2020) using LM-EVAL (Gao et al., 2023). The average of WinoGrande, PIQA, and ARC-easy (accuracy), and HellaSwag and ARC-challenge (length-normalized accuracy) is reported for experiments. More evaluations are placed in Appendix B.
@@
Our kernels are adapted from the 4-bit kernels (Frantar et al., 2024) and RoPE kernels (Hsu et al., 2025). All computations are in BF16, except for correlation matrices for structured sorting and Cholesky decomposition for GPTQ are calculated in FP32. Detailed parameters are placed at Appendix G. The weight-sorting, masked fine-tuning, and quantization are computed on an A6000 GPU with 48GB memory in one-shot for enabling adaptive pruning on the device. We profile the latency on A6000 and Orin Nano 8GB, as our experimental cloud and edge platforms. We report the average latency of twenty profiles after five warmup runs.
Table 4: (One-pass adaptive pruning.) We evaluate UniQL in one run across pruning rates. R.size stands for reduction of model size ($\times$). UniQL enables all pruning rates in single pass. The symbols $\diamond$, $\dagger$, and $\ddagger$ denote Transformers, Mamba-Transformer hybrids, and Mamba models, respectively. represents the FP16 embeddings and output layers as per the official implementation. We apply GPTQ (Frantar et al., 2023) on MoDeGPT (Lin et al., 2025) denoted as .
@@ 4.3 Compression Time
In Table 6, we compare the compression time of UniQL against MoDeGPT (Lin et al., 2025), noted for being training-free, and SVD-LLM (Wang et al., 2025b), which involves fine-tuning, both state-of-the-art algorithms for Transformer compression. Our matrix decomposition is $22\times$ (0h19m vs. 7h03m) faster than MoDeGPT and $1.8\times$ (0h19m vs. 0h35m) faster than SVD-LLM, as UniQL avoids pseudo-inverse and SVD decomposition for large MLP weight matrices. With masked fine-tuning (FT), UniQL remains quicker (6h59m) than both MoDeGPT (7h03m) and SVD-LLM (15h57m). MoDeGPT suffers from the high computation cost of performing pseudo-inverse on large weight matrices. SVD-LLM splits weights into two successive layers, $\mathbf{U}$ and $\mathbf{V}$, and carries out independent fine-tuning for each, leading to a longer fine-tuning duration. Lastly, post-training quantization (PTQ) takes an extra forty minutes. Our compression algorithm is one-time $O(1)$ with respect to the number of compression rates, compared to $O(n)$ for MoDeGPT and SVD-LLM.
@@ 4.4 Model Size and Latency Profiling
We evaluate the model size and latency of UniQL and compare them with AWQ (Lin et al., 2024a) from TensorRT-MO11footnotemark: 1 (NVIDIA, 2024, 2023) (TRT-AWQ) and HQQ (Badri and Shaji, 2023) in TorchAO11footnotemark: 1 (torchao, 2024) (TAO-HQQ). Both libraries are weight-only (i.e., W4A16) quantization frameworks in production. We show the model size in Table 6. UniQL quantizes models in head-to-toe fashion, i.e., from embeddings, backbone layers, and output heads all at 4-bit, resulting in smaller model size (4.1GB vs. 5.7GB) compared to TRT-AWQ and TAO-HQQ with minimal accuracy drops. Our model enables all compression rates and on-device structured pruning, providing an elastic $3.9\times$$5.7\times$ memory reductions. We profile the time-per-output-token (TPOT, i.e., generation) and time-to-last-token (TTLT, i.e., prefilling and generation) on A6000 and Nano 8G, and show $2.7\times$$3.4\times$ throughput improvements in generation, outperforming TRT-AWQ and TAO-HQQ, as shown in Table 8 and Table 8. On the Nano 8G, our model is $1.7\times$ faster than TAO-HQQ in TPOT. By pruning 35% of weights in our 4-bit model, our models generate $2.1\times$ faster than TAO-HQQ.
@@ Acknowledgments
This work was supported in part by NSF CCF Grant No. 2107085, iMAGiNE - the Intelligent Machine Engineering Consortium at UT Austin, UT Cockrell School of Engineering Doctoral Fellowships, NSF CAREER Grant No. 2339084, Nvidia research gift, and Taiwan’s NSTC Grant No. 111-2221-E-A49-148-MY3.
@@ B.1 Comparison with Additional Baselines
we include more common baselines in Table 11, with all methods evaluated in FP16 to ensure a fair and controlled setting. We follow the experimental setup used in MoDeGPT [Lin et al., 2025] and append our Llama-3.1-8B results at the 25% pruning rate to those reported in their work. The numbers, except for UniQL, are transcribed from MoDeGPT. The results show that UniQL consistently outperforms prior pruning-based methods at the 25% compression level, and UniQL-ft further boosts accuracy, achieving the strongest performance across all evaluated tasks.
@@ B.2 Evaluation on the MMLU Dataset
We test Llama-3.1-8B, Qwen-2.5-7B, and Nemotron-H-8B, and report five-shot accuracy on the MMLU dataset [Hendrycks et al., 2020] with a batch size of eight. The MMLU dataset is a large multitasking dataset, covering 57 subjects of varying difficulty. Our pruned models maintain competitive accuracy on the challenging dataset and outperform MoDeGPT [Lin et al., 2025] and SVD-LLM [Wang et al., 2025b]. Nemotron-H, the Mamba-Transformer hybrid model, shows a large accuracy drop, but recovered with our low-cost masked fine-tuning. We compare our method with AWQ [Lin et al., 2024a] implemented in the TensorRT framework [NVIDIA, 2023, 2024] (TRT-AWQ) and HQQ [Badri and Shaji, 2023] in TorchAO [torchao, 2024] (TAO-HQQ). Our 4-bit models perform comparably to these state-of-the-art PTQ frameworks while offering broader model support.
Table 12: (Five-shot MMLU.) We compare UniQL against baselines under different settings. represent the FP16 embeddings and output layers as per the official implementation. denotes the GPTQ [Frantar et al., 2023] implemented on all models as an additional baseline.
Method +FT +PTQ W-bit Prun. p% Llama-3.1-8B Qwen-2.5-7B Nemotron-H-8B
@@ B.3 Evaluation on Coding Tasks
We present the evaluation results on the MBPP+ [Austin et al., 2021, Gao et al., 2023] coding benchmark in Table 13. We apply UniQL to Llama-3.1-8B-Instruct and compare its performance against SVD-LLM [Wang et al., 2025b] and MoDeGPT [Lin et al., 2025] under various pruning ratios and bit-width configurations. The MBPP+ results obtained under batch size 1 and 0-shot settings following common practice. These results demonstrate UniQL’s ability to maintain competitive performance compared to SVD-LLM and MoDeGPT while reducing model size.